# Fixation and extinction in time-fluctuating spatially structured metapopulations

### v2

@@ IV.2 Strong bottlenecks: $\psi(m,K_{-})<1$
Figure 5: Typical single realisations of $N/\Omega$ (black), $N_{M}/\Omega$ (red), $N_{W}/\Omega$ (blue), and $K(t)$ (grey) vs time for different values of $K_{-}$ and $\nu$. (a,b): Here, $\nu=10^{-4}$ and $K_{-}=8$. In (a), $K=K_{-}$ at $t=0$ and $M$ and then $W$ quickly go extinct. In (b), $K=K_{+}$ at $t=0$ and $M$ fixes the population while $W$ goes extinct. (c,d): Here, $\nu=10^{-2}$ and $K_{-}=8$. In (c), mutants survive the first few bottlenecks but their abundance is low leading to the fixation of $W$ and removal of $M$ after four bottlenecks ($t\gtrsim 1000$). In (d), mutants survive the first bottlenecks and spread in the mild state where they recolonise and invade demes. They are eventually able to fix the population. (e,f): Here, $\nu=10$, and $K_{-}=4$ in (e) and $K_{-}=10$ in (f). $K(t)$ switches very frequently and is not shown for clarity. In (e), the deme size is $n\approx 2K_{-}=8$ and the dynamics is dominated by deme extinction leading to the rapid extinction of the metapopulation. In (e), the deme size is $n\approx 2K_{-}=20$ and there is $M/W$ competition that leads to fixation of $M$ and extinction of $W$ after a typical time $t\sim\theta^{\rm clique}(2K_{-})\gtrsim 10^{4}$ (not shown). Similar results are obtained on other regular graphs $\rm G$}; see text. Other parameters are $\Omega=10$, $s=0.1$, $m=10^{-4}$, and $K_{+}=200$. In all panels, initially there is a single $M$ deme and $\Omega-1$ demes occupied by $W$.
@@
Figure 6: Near-optimal condition for the idealised treatment strategy. $(\nu,K_{-})$ heatmaps of $\Phi$, $\Theta$, $\Theta_{E}$ and $(1-\Phi)(1-\Theta/\Theta_{E})$ for a clique metapopulation. The migration rate is $m=10^{-3}$ in (a-d) and $m=10^{-4}$ in (e-h). White space in panels (c) and (g) indicates where at least one realisation for those parameters did not reach extinction by $t=10^{5}$, i.e. $\Theta_{E}\sim 10^{5}$ or larger. Grey lines in panels (d) and (h) show the near-optimal conditions for the idealised treatment strategy: $\psi(m,K_{-})<1$ below the top horizontal line, $mK_{+}\theta_{E}>1$ above the bottom horizontal line, and $\nu\theta_{E}>1$ above the curved line, while the vertical line indicates where $\nu<1$ and $\theta_{E}$ from Eq. (15). The near-optimal treatment conditions is the yellowish cloud at the centre of the area enclosed by these lines. Similar results are obtained on other regular graphs $\rm G$}; see text and Fig. 10. Other parameters $\Omega=16$, $s=0.1$, and $K_{+}=200$. In all panels, initially there is a single $M$ deme and $\Omega-1$ demes occupied by $W$.
@@
When $\psi(m,K_{-})<1$, each deme can undergo strong bottlenecks; see Fig. 3(a) and below. In the harsh environment $\xi=-1$ ($K=K_{-}$), the entire metapopulation experiences extinction, in an observable time $\theta_{E}(K_{-},\Omega)$, denoted $\theta_{E}$ here for conciseness. However, in the mild state $\xi=1$ ($K=K_{+}$), deme extinction can be neglected and each site can be regarded as being occupied by either $W$ or $M$ cells. The dynamics in the harsh state is thus dominated by extinction, whereas the $M/W$ competition characterising the mild state is aptly captured by the two-state coarse-grained description of Section III.1. Environmental switching thus couples regimes that are dominated in turn by deme extinction and $M/W$ competition, yielding complex dynamical scenarios whose analysis is difficult. However, we can gain valuable insight by considering first the limits $\nu\ll 1$, $\nu\gg 1$, and then the case of intermediate switching where $\nu\lesssim 1$.
@@
When the environment varies very slowly, $\nu\ll 1$, the $K(t)$ remains at its initial value for long periods, that is $K=K_{\pm}$ if $\xi(0)=\pm 1$ each with a probability $1/2$. On the one hand, if initially $\xi=-1$ (harsh environment), $K=K_{-}$ and each deme is prone to extinction after a mean time $\tau_{E}(K_{-})$, which eventually leads to the collapse of the metapopulation after a mMET $\theta_{E}\approx e^{K_{-}}\ln{(\Omega)}/K_{-}$ when $\Omega\gg 1$ and $K_{-}\gg 1$; see Eqs. (7) and (15) and Fig. 5(a). On the other hand, if initially $\xi=1$ (mild condition), $n\approx K_{+}$ and there is $M/W$ competition characterised by the fixation of $M$ with a probability $\phi(K_{+})$ approaching 1 when $K_{+}s\gg 1$; see Eq. (14) and Fig. 5(b). As a result, when $\nu\ll 1$ and $K_{+}s\gg 1$, there are two equally likely outcomes illustrated in Fig. 5: either the extinction of the metapopulation in a mean time $\theta_{E}$ as in Fig. 5(a), or the fixation of $M$ after a mean time $\theta^{\rm G}(K_{+})$ as in Fig. 5(b).
@@ A.1 Master equation
As discussed in Sec. II, the individual-based model is a continuous-time multivariate Markov process defined by the reaction and transition rates Eq. (1)-(4). The intra and inter-deme dynamics is encoded in a master equation for the probability $P(\{n_{W},n_{M}\},\xi,t)$ that at time $t$ the metapopulation is in the environmental state $\xi$ and configuration $\{n_{W},n_{M}\}\equiv(\dots,n_{W}(x),n_{M}(x),\dots)$, where $n_{W/M}(x)$ denotes the number of cells of type $W/M$ in deme $x=1,\dots,\Omega$. The master equation (ME) for the metapopulation dynamics subject to environmental switching on a regular graph ${\rm G}=\{{\rm clique,cycle,grid}\}$ with degrees (or number of nearest neighbours) given by $q_{\text{clique}}=\Omega-1$, $q_{\text{cycle}}=2$, or $q_{\text{grid}}=4$ reads:
@@
While the ME Eq. (28) holds for any regular graphs $\rm G$}, in our examples we consider specifically the regular circulation graphs ${\rm G}=\{{\rm clique,cycle,grid}\}$. The space-dependent individual-based dynamics encoded in the ME Eq. (28) has been simulated using the Monte Carlo method described in Appendix I. It is worth noting that demographic fluctuations eventually lead to the extinction of the entire metapopulation, in all regimes. However, this phenomenon occurs after a time growing dramatically with the system size, and it can generally not be observed in sufficiently large metapopulations; see Fig. 2(f).
@@ A.2 Eco-evolutionary dynamics of a single deme
For an isolated deme, there is only intra-deme dynamics according to the birth-death process defined by Eqs. (1) and (2). The ME for the dynamics of a single isolated deme is thus given by setting $\Omega=1$ and $m=0$ in Eq. (28); see Appendix A.1. The corresponding intra-deme dynamics in a time-varying environment can be simulated using exact methods like the Gillespie algorithm [97], as in Refs [16, 17, 18, 23, 20, 21, 22, 24]. It is instructive to ignore all forms of fluctuations and consider the mean-field approximation of an isolated deme dynamics subject to a constant carrying capacity $K\gg 1$. Following Refs. [16, 17], with the transition rates Eq. (2), the mean-field eco-evolutionary dynamics of a single isolated deme is characterised by rate equations for the size $n$ of the deme and the fraction $x=n_{M}/n$ of mutants in the deme, which read:
@@
When demographic fluctuations can be neglected and randomness only arises from environmental variability via Eq. (6), the intra-deme dynamics of an isolated deme is well-captured by the piecewise deterministic Markov process (PDMP) for $n$, obtained by ignoring demographic fluctuations [85, 16, 17]. In the realm of the PDMP approximation, the deme size thus satisfies a deterministic logistic equation in each environmental state, subject to a carrying capacity that switches when the environment changes ($K=K_{\pm}$ when $\xi=\pm 1$), yielding the following $n$-PDMP [16, 17, 18, 22, 32, 21]:
$\dot{n}=\begin{cases}n\left(1-\frac{n}{K_{-}}\right)&\text{if }\xi=-1,\\ n\left(1-\frac{n}{K_{+}}\right)&\text{if }\xi=1,\end{cases}$ (30)
@@
Figure 7: Quasi-stationary probability density of deme population size ($n$-QPSD) distribution on clique metapopulations for various parameters and its $n$-PDMP approximation given by Eqs. (19) and (31). Red and blue bars show data for the $n$-QPSD conditioned on $K(t)=K_{-}$ and $K(t)=K_{+}$, respectively. Red, blue, and black solid lines are the $n$-PDMP stationary densities $p_{-}(n;\nu)$, $p_{+}(n;\nu)$, and $p(n;\nu)$, respectively, given by Eqs. (19) and (31). (a-c) are for $m=10^{-3}$, (d-f) for $m=10^{-2}$, and (g-i) for $m=10^{-1}$. We have $\nu=10^{-3}$ in (a,d,g), $\nu=10^{-1}$ in (b,e,h), and $\nu=10^{2}$ in (c,f,i). Other parameters are $\Omega=16$, $K_{+}=200$, and $K_{-}=20$. All represent a single realisation tracked until $t=10^{5}$.
@@ Appendix C Deme and metapopulation mean extinction times
In this section, we discuss the process of extinction of a single deme that has a constant carrying capacity $K$ when $\psi(m,K)\ll 1$; see Sec. III.2. In this extinction-dominated regime, we can assume that the deme size rapidly fluctuates about $K$, and extinction occurs from a deme of constant size $n(x)=K$ prior to any invasion. Without loss of generality (see below), we hence assume that extinction occurs from entirely occupied demes, with the metapopulation consisting only of $W$ and $M$ demes, all of size $K$. In this representation, the dynamics of a $W/M$ deme when $\psi\ll 1$ is given by the birth-death process of Sec. II with transition rates $T_{W/M}^{+}=n_{W/M}$ and $T_{W/M}^{-}=n_{W/M}^{2}/K$, subject to an absorbing boundary at $n_{W/M}=0$; see Eq. (2). Clearly therefore, the deme dynamics is independent of its type, and the deme mean extinction time (dMET) is the average time to reach $n_{W/M}=0$ and is the same for $W$ and $M$ demes (dMET is independent of $s$). A classical calculation (see, e.g. Sec. 6.7 in Ref. [72]) yields the following expression for an initially fully occupied deme of size $K$:
@@
In the vein of Ref. [59] the recolonisation-birth rate of occupied demes is $B(j)=mKj(1-j/\Omega)$, corresponding to a logistic growth with a rate proportional to the expected number of migrations from an occupied deme $mK$. Here, the extinction rate is $D(j)=j/\tau_{E}$ and is inversely proportional to the mean local extinction time, that is the dMET. With $\psi=mK\tau_{E}$, using $\prod_{l=n}^{j-1}(1-\frac{l}{\Omega})=\frac{1}{\Omega^{j-n}}\frac{(\Omega-n)!}{(\Omega-j)!}$, Eq. (38) can be rewritten as
@@ Appendix E Average number of active edges on the square grid and the influence of the spatial structure
In Figs. 2(a,c) and 4(c), we have found that the spatial structure has a barely noticeable influence on the fixation probability and mean fixation time when the carrying capacity is constant and in time-varying environments in the regime of weak bottlenecks. Fig. 10 shows the heatmaps on a cycle and a grid metapopulation for the “idealised treatment strategy” proposed in Sec. IV.2, which are almost identical. This is in accord with Eqs. (27) predicting that the same migration rate yields the same near optimal conditions for the heatmaps of metapopulation on any regular graph, here a cycle and a grid. Simulation results confirm spatial structure is only responsible for minor quantitative changes in the region of the heatmaps corresponding to the near-optimal “treatment conditions”. This stems from the removal scenario characterising the idealised treatment strategy being due to deme extinction which is mostly independent of $\rm G$} and $m$.
Appendix F Further details on the intermediate dynamics in static environments
@@
Figure 11: $\phi_{\rm int}$ vs. $K$ for $s=0.1$ (blue) and $s=0.01$ (red) on clique (solid lines / crosses), cycle (dashed lines / circles), and a grid (dotted lines / triangles). The different symbols and lines are almost indistinguishable. $\theta_{\rm int}$ vs. $K$ for the same parameters for a clique metapopulation. The vertical solid and dashed line indicate where the number of occupied demes $\Omega_{\text{occ}}\in[1,\Omega-1]$; see text. The dotted line indicates where $\psi=\Omega K$, i.e. every individual migrates in the time required for an independent deme extinction. Markers show simulation results, solid lines are predictions of Eq. (44), and dash-dot lines are predictions of Eq. (14). Other parameters are $m=10^{-4}$ and $\Omega=16$.
@@
In this section, we provide further details about the analysis of the intermediate dynamics in static environments; see Sec. III.3. Here, we complete the result given in the main text by considering metapopulation intermediate dynamics on a regular graph $\rm G$} ($\rm G$}-graph), obtaining the explicit results $\phi_{\rm int}^{\rm G}$ and $\theta_{\rm int}^{\rm G}$ for ${\rm G}=\{{\rm clique,cycle,grid}\}$ reported in Fig. 11 when the metapopulation initially consists of a single $M$ deme and $\Omega-1$ demes occupied by $W$. This region is concretely defined for values of $\psi$ such that $1<\psi<K\Omega$ for $\Omega\gg 1$, where the lower bound ensures $\Omega_{\text{occ}}=1$ from Eq. (9) and the upper bound is obtained by considering the limiting case where every individual is expected to migrate on the timescale of deme extinction. Using that $\psi\approx me^{K}$ for $K\gg 1$, we obtain the condition $\ln(1/m)\lesssim K\lesssim\ln(\Omega K/m)$.
@@
The intermediate regime is characterised by $M/W$ competition and deme extinction. Therefore, in addition to invasions, a $W$ deme may become an $M$ deme through extinction followed by a recolonisation, i.e. $W\rightarrow\emptyset\rightarrow M$, where $\emptyset$ indicates an extinct deme. Similarly, an $M$-deme can be changed into a $W$-deme via $M\rightarrow\emptyset\rightarrow W$. We assume that there is initially a single $M$ deme in the metapopulation (and $\Omega-1$ demes of type $W$). With a probability $p_{\text{surv}}$ (see below), the initial $M$ deme survives the short transient as the metapopulation reaches the dynamical equilibrium that consists of the $M$ deme, the remaining $W$ demes, and extinct demes. The number of $M$ demes $i=0,1,\dots,\Omega_{\rm occ}$ grows and shrinks through invasions and extinction-recolonisation events. We assume that immediately $\Omega_{E}=\Omega/\psi$ demes go extinct, so that the metapopulation quickly reaches its equilibrium occupancy $\Omega_{\text{occ}}=\Omega\left(1-1/\psi\right)$. In this dynamical equilibrium, a $W$ deme can become an $M$ deme via $W\rightarrow\emptyset$ ($W$ deme extinction) at rate $r_{{\rm ext},W}$ followed by $\emptyset\rightarrow M$ (recolonisation by $M$) at rate $r_{{\rm rec},M}^{\rm G}$. The overall extinction-recolonisation reaction $W\rightarrow\emptyset\rightarrow M$ thus occurs at rate $1/(1/r_{{\rm ext},W}+1/r_{{\rm rec},M}^{\rm G})$. Here, the rate of $W$ deme extinction is $r_{{\rm ext},W}=(\Omega_{\text{occ}}-i)/\tau_{E}$ and $\tau_{E}$ is given by Eq. (7). We proceed similarly for the extinction of an $M$ deme and its recolonisation into a $W$ site according to $M\rightarrow\emptyset\rightarrow W$. Taking also into account the rate of invasion (see Eq. (11)) the number of $M$ demes $i$ on a $\rm G$}-regular graph varies according to the transition rates
@@
It is important to note that the $\widetilde{E}_{\rm G}(i)$ in these expressions explicitly represent the number of active edges between $M$ and $W$ on the metapopulation given $i$ mutant demes, taking into account the presence of extinct demes. Therefore, their expressions for a given graph structure $\rm G$ generally differ from those considered in the competition-dominated regime, denoted $E_{\rm G}(i)$. With these rates, we can solve the following first-step analysis equations for the probability $\phi_{{\rm int},i}^{\rm G}$ that the dynamical equilibrium comprising initially $i$ demes of type $M$ consists of occupied $M$ demes and extinct demes after a mean time $\theta_{{\rm int},i}^{\rm G}$:

### v3

@@ IV.2 Strong bottlenecks: $\psi(m,K_{-})<1$
Figure 5: Typical single realisations of $N/\Omega$ (black), $N_{M}/\Omega$ (red), $N_{W}/\Omega$ (blue), and $K(t)$ (grey) against time for different values of $K_{-}$ and $\nu$. (a,b): Here, $\nu=10^{-4}$ and $K_{-}=8$. In (a), $K=K_{-}$ at $t=0$ and $M$ and then $W$ quickly go extinct. In (b), $K=K_{+}$ at $t=0$ and $M$ fixes the population while $W$ goes extinct. (c,d): Here, $\nu=10^{-2}$ and $K_{-}=8$. In (c), mutants survive the first few bottlenecks but their abundance is low leading to the fixation of $W$ and removal of $M$ after four bottlenecks ($t\gtrsim 1000$). In (d), mutants survive the first bottlenecks and spread in the mild state where they recolonise and invade demes. They are eventually able to fix the population. (e,f): Here, $\nu=10$, and $K_{-}=4$ in (e) and $K_{-}=10$ in (f). $K(t)$ switches very frequently and is not shown for clarity. In (e), the deme size is $n\approx 2K_{-}=8$ and the dynamics is dominated by deme extinction leading to the rapid extinction of the metapopulation. In (e), the deme size is $n\approx 2K_{-}=20$ and there is $M/W$ competition that leads to fixation of $M$ and extinction of $W$ after a typical time $t\sim\theta^{\rm clique}(2K_{-})\gtrsim 10^{4}$ (not shown). Similar results are obtained on other regular graphs $\rm G$}; see text. Other parameters are $\Omega=10$, $s=0.1$, $m=10^{-4}$, and $K_{+}=200$. In all panels, initially there is a single $M$ deme and $\Omega-1$ demes occupied by $W$.
@@
Figure 6: Near-optimal condition for the idealised treatment strategy. $(\nu,K_{-})$ heatmaps of $\Phi$, $\Theta$, $\Theta_{E}$ and $(1-\Phi)(1-\Theta/\Theta_{E})$ for a clique metapopulation. The migration rate is $m=10^{-3}$ in (a-d) and $m=10^{-4}$ in (e-h). White space in panels (c) and (g) indicates where at least one realisation for those parameters did not reach extinction by $t=10^{5}$, i.e. $\Theta_{E}\sim 10^{5}$ or larger. Grey lines in panels (d) and (h) are show the near-optimal conditions for the idealised treatment strategy, given by Eq. (27): $\psi(m,K_{-})<1$ below the top horizontal line, $mK_{+}\theta_{E}>1$ above the bottom horizontal line, and $\nu\theta_{E}>1$ above the curved line, while the vertical line indicates where $\nu<1$. Here, $\theta_{E}$ is obtained from Eq. (15). The near-optimal treatment conditions is the yellowish cloud at the centre of the area enclosed by these lines. Similar results are obtained on other regular graphs $\rm G$}; see text and Fig. 10. Other parameters $\Omega=16$, $s=0.1$, and $K_{+}=200$. In all panels, initially there is a single $M$ deme and $\Omega-1$ demes occupied by $W$.
@@
When $\psi(m,K_{-})<1$, each deme can undergo strong bottlenecks; see Fig. 3(a) and below. In the harsh environment $\xi=-1$ ($K=K_{-}$), the entire metapopulation experiences extinction, in an observable time $\theta_{E}(K_{-},\Omega)$, denoted $\theta_{E}$ here for conciseness. However, in the mild state $\xi=+1$ ($K=K_{+}$), deme extinction can be neglected and each site can be regarded as being occupied by either $W$ or $M$ cells. The dynamics in the harsh state is thus dominated by extinction, whereas the $M/W$ competition characterising the mild state is aptly captured by the two-state coarse-grained description of Section III.1. Environmental switching thus couples regimes that are dominated in turn by deme extinction and $M/W$ competition, yielding complex dynamical scenarios whose analysis is difficult. However, we can gain valuable insight by considering first the limits $\nu\ll 1$, $\nu\gg 1$, and then the case of intermediate switching where $\nu\lesssim 1$.
@@
When the environment varies very slowly, $\nu\ll 1$, the $K(t)$ remains at its initial value for long periods, that is $K=K_{\pm}$ if $\xi(0)=\pm 1$ each with a probability $1/2$. On the one hand, if initially $\xi=-1$ (harsh environment), $K=K_{-}$ and each deme is prone to extinction after a mean time $\tau_{E}(K_{-})$, which eventually leads to the collapse of the metapopulation after a mMET $\theta_{E}\approx e^{K_{-}}\ln{(\Omega)}/K_{-}$ when $\Omega\gg 1$ and $K_{-}\gg 1$; see Eqs. (7) and (15) and Fig. 5(a). On the other hand, if initially $\xi=+1$ (mild condition), $n\approx K_{+}$ and there is $M/W$ competition characterised by the fixation of $M$ with a probability $\phi(K_{+})$ approaching 1 when $K_{+}s\gg 1$; see Eq. (14) and Fig. 5(b). As a result, when $\nu\ll 1$ and $K_{+}s\gg 1$, there are two equally likely outcomes illustrated in Fig. 5: either the extinction of the metapopulation in a mean time $\theta_{E}$ as in Fig. 5(a), or the fixation of $M$ after a mean time $\theta^{\rm G}(K_{+})$ as in Fig. 5(b).
@@ A.1 Master equation
As discussed in Sec. II, the individual-based model is a continuous-time multivariate Markov process defined by the reaction and transition rates Eq. (1)-(4). The intra and inter-deme dynamics is encoded in a master equation for the probability $P(\{n_{W},n_{M}\},\xi,t)$ that at time $t$ the metapopulation is in the environmental state $\xi$ and configuration $\{n_{W},n_{M}\}\equiv(\dots,n_{W}(x),n_{M}(x),\dots)$, where $n_{W/M}(x)$ denotes the number of cells of type $W/M$ in deme $x=1,\dots,\Omega$. The master equation (ME) for the metapopulation dynamics subject to environmental switching on a regular graph ${\rm G}=\{{\rm clique,cycle,grid}\}$ with degrees (or number of nearest neighbours) given by $q_{\text{clique}}=\Omega-1$, $q_{\text{cycle}}=2$, or $q_{\text{grid}}=4$ reads
@@
While the ME Eq. (28) holds for any regular graphs $\rm G$}, in our examples we consider specifically the regular circulation graphs ${\rm G}=\{{\rm clique,cycle,grid}\}$. The space-dependent individual-based dynamics encoded in the ME Eq. (28) has been simulated using the Monte Carlo method described in Appendix I. It is worth noting that demographic fluctuations eventually lead to the extinction of the entire metapopulation, in all regimes. However, this phenomenon occurs after a time growing dramatically with the system size, and it can generally not be observed in sufficiently large metapopulations; see Fig. 2(b,bottom right).
@@ A.2 Eco-evolutionary dynamics of a single deme
For an isolated deme, there is only intra-deme dynamics according to the birth-death process defined by Eqs. (1) and (2). The ME for the dynamics of a single isolated deme is thus given by setting $\Omega=1$ and $m=0$ in Eq. (28); see Appendix A.1. The corresponding intra-deme dynamics in a time-varying environment can be simulated using exact methods like the Gillespie algorithm [97], as in Refs [16, 17, 18, 23, 20, 21, 22, 24]. It is instructive to ignore all forms of fluctuations and consider the mean-field approximation of an isolated deme dynamics subject to a constant carrying capacity $K\gg 1$. Following Refs. [16, 17], with the transition rates Eq. (2), the mean-field eco-evolutionary dynamics of a single isolated deme is characterised by rate equations for the size $n$ of the deme and the fraction $x=n_{M}/n$ of mutants in the deme, which read
@@
When demographic fluctuations can be neglected and randomness only arises from environmental variability via Eq. (6), the intra-deme dynamics of an isolated deme is well-captured by the piecewise deterministic Markov process (PDMP) for $n$, obtained by ignoring demographic fluctuations [85, 16, 17]. In the realm of the PDMP approximation, the deme size thus satisfies a deterministic logistic equation in each environmental state, subject to a carrying capacity that switches when the environment changes ($K=K_{\pm}$ when $\xi=\pm 1$), yielding the $n$-PDMP [16, 17, 18, 22, 32, 21]
$\dot{n}=\begin{cases}n\left(1-\frac{n}{K_{-}}\right)&\text{if }\xi=-1,\\ n\left(1-\frac{n}{K_{+}}\right)&\text{if }\xi=+1,\end{cases}$ (30)
@@
Figure 7: Quasi-stationary probability density of deme population size ($n$-QPSD) distribution on clique metapopulations for various parameters and its $n$-PDMP approximation given by Eqs. (19) and (31). Red and blue bars show data for the $n$-QPSD conditioned on $K(t)=K_{-}$ and $K(t)=K_{+}$, respectively. Red, blue, and black solid lines are the $n$-PDMP stationary densities $p_{-}(n;\nu)$, $p_{+}(n;\nu)$, and $p(n;\nu)$, respectively, given by Eqs. (19) and (31). Panels (a-c) are for $m=10^{-3}$, (d-f) for $m=10^{-2}$, and (g-i) for $m=10^{-1}$. We have $\nu=10^{-3}$ in (a,d,g), $\nu=10^{-1}$ in (b,e,h), and $\nu=10^{2}$ in (c,f,i). Other parameters are $\Omega=16$, $K_{+}=200$, and $K_{-}=20$. All represent a single realisation tracked until $t=10^{5}$.
@@ Appendix C Deme and metapopulation mean extinction times
In this section, we discuss the process of extinction of a single deme that has a constant carrying capacity $K$ when $\psi(m,K)\ll 1$; see Sec. III.2. In this extinction-dominated regime, we can assume that the deme size rapidly fluctuates about $K$, and extinction occurs from a deme of constant size $n(x)=K$ prior to any invasion. Without loss of generality (see below), we hence assume that extinction occurs from entirely occupied demes, with the metapopulation consisting only of $W$ and $M$ demes, all of size $K$. In this representation, the dynamics of a $W/M$ deme when $\psi\ll 1$ is given by the birth-death process of Sec. II with transition rates $T_{W/M}^{+}=n_{W/M}$ and $T_{W/M}^{-}=n_{W/M}^{2}/K$, subject to an absorbing boundary at $n_{W/M}=0$; see Eq. (2). Clearly therefore, the deme dynamics is independent of its type, and the deme mean extinction time (dMET) is the average time to reach $n_{W/M}=0$ and is the same for $W$ and $M$ demes (dMET is independent of $s$). A classical calculation (see, e.g. Sec. 6.7 in Ref. [72]) for an initially fully occupied deme of size $K$ yields
@@
In the vein of Ref. [59], the recolonisation-birth rate of occupied demes is $B(j)=mKj(1-j/\Omega)$, corresponding to a logistic growth with a rate proportional to the expected number of migrations from an occupied deme $mK$. Here, the extinction rate is $D(j)=j/\tau_{E}$ and is inversely proportional to the mean local extinction time, that is the dMET. With $\psi=mK\tau_{E}$, using $\prod_{l=n}^{j-1}(1-\frac{l}{\Omega})=\frac{1}{\Omega^{j-n}}\frac{(\Omega-n)!}{(\Omega-j)!}$, Eq. (38) can be rewritten as
@@ Appendix E Average number of active edges on the square grid and the influence of the spatial structure
In Figs. 2(a,top) and 4(c), we have found that the spatial structure has a barely noticeable influence on the fixation probability under both constant carrying capacity and in the regime of weak bottlenecks. The mean fixation time also appears unchanged when the carrying capacity is constant. Fig. 10 shows the heatmaps on a cycle and a grid metapopulation for the “idealised treatment strategy” proposed in Sec. IV.2, which are almost identical. This is in accord with Eqs. (27) predicting that the same migration rate yields the same near optimal conditions for the heatmaps of metapopulation on any regular graph, here a cycle and a grid. Simulation results confirm spatial structure is only responsible for minor quantitative changes in the region of the heatmaps corresponding to the near-optimal “treatment conditions”. This stems from the removal scenario characterising the idealised treatment strategy being due to deme extinction which is mostly independent of $\rm G$} and $m$.
Appendix F Intermediate dynamics in static environments
@@
Figure 11: $\phi_{\rm int}$ vs. $K$ for $s=0.1$ (blue) and $s=0.01$ (red) on clique (solid lines / crosses), cycle (dashed lines / circles), and a grid (dotted lines / triangles). The different symbols and lines are almost indistinguishable. $\theta_{\rm int}$ vs. $K$ for the same parameters for a clique metapopulation. The vertical solid and dashed line indicate where the number of occupied demes $\Omega_{\text{occ}}\in[1,\Omega-1]$; see text. The dotted line indicates where $\psi=K\Omega$, i.e. every individual migrates in the time required for an independent deme extinction. Markers show simulation results, solid lines are predictions of Eq. (44), and dash-dot lines are predictions of Eq. (14). Other parameters are $m=10^{-4}$ and $\Omega=16$.
@@
When $\psi\gtrsim 1$ with $mK<1$ (slow migration) in static environments, migration and deme extinction occur on the timescale $\tau_{E}$. In the long run, deme recolonisations and extinctions balance each other, yielding a dynamical equilibrium consisting of $\Omega_{\rm occ}=\Omega(1-1/\psi)$ occupied demes and $\Omega-\Omega_{\rm occ}$ empty demes; see Fig. 9. In this regime, the three-state coarse-grained description of the dynamical equilibrium consists of a random mixture of empty demes, and occupied $W$ and $M$ demes; see Fig. 1(d). After a mean time $\theta_{\rm int}^{\rm G}$ the metapopulation reaches the dynamical equilibrium consisting of a fraction $1-\Omega_{\rm occ}/\Omega=1/\psi$ empty demes, and the remaining demes are all of either type $M$ or $W$ with probability $\phi^{G}_{\rm int}$ and $1-\phi_{\rm int}^{\rm G}$, respectively. The dynamical equilibrium is thus defined by the quantities $\psi$, given by Eq. (8), and $\phi_{\rm int}^{\rm G}$ which is the probability that mutants $M$ take over the $\Omega_{\rm occ}=\Omega(1-1/\psi)$ occupied demes, where the unconditional mean fixation time is given by $\theta_{\rm int}^{\rm G}$. In this section, we complete the results given in the main text by considering metapopulation intermediate dynamics on a regular graph $\rm G$}, obtaining the explicit results $\phi_{\rm int}^{\rm G}$ and $\theta_{\rm int}^{\rm G}$ for ${\rm G}=\{{\rm clique,cycle,grid}\}$ reported in Fig. 11 when the metapopulation initially consists of a single $M$ deme and $\Omega-1$ demes occupied by $W$. This region is concretely defined for values of $\psi$ such that $1<\psi<K\Omega$ for $\Omega\gg 1$, where the lower bound ensures $\Omega_{\text{occ}}=1$ from Eq. (9) and the upper bound is obtained by considering the limiting case where every individual is expected to migrate on the timescale of deme extinction. Using that $\psi\approx me^{K}$ for $K\gg 1$, we obtain the condition $\ln(1/m)\lesssim K\lesssim\ln(K\Omega/m)$. These bounds are illustrated by the vertical lines in Fig. 11.
@@
The intermediate regime is characterised by $M/W$ competition and deme extinction. Therefore, in addition to invasions, a $W$ deme may become an $M$ deme through extinction followed by a recolonisation, i.e. $W\rightarrow\emptyset\rightarrow M$, where $\emptyset$ indicates an extinct deme. Similarly, an $M$-deme can be changed into a $W$-deme via $M\rightarrow\emptyset\rightarrow W$. We assume that there is initially a single $M$ deme in the metapopulation (and $\Omega-1$ demes of type $W$). With a probability $p_{\text{surv}}$ (see below), the initial $M$ deme survives the short transient as the metapopulation reaches the dynamical equilibrium that consists of the $M$ deme, the remaining $W$ demes, and extinct demes. The number of $M$ demes $i=0,1,\dots,\Omega_{\rm occ}$ grows and shrinks through invasions and extinction-recolonisation events. We assume that immediately $\Omega_{E}=\Omega/\psi$ demes go extinct, so that the metapopulation quickly reaches its equilibrium occupancy $\Omega_{\text{occ}}=\Omega\left(1-1/\psi\right)$. In this dynamical equilibrium, a $W$ deme can become an $M$ deme via $W\rightarrow\emptyset$ ($W$ deme extinction) at rate $r_{{\rm ext},W}$ followed by $\emptyset\rightarrow M$ (recolonisation by $M$) at rate $r_{{\rm rec},M}^{\rm G}$. The overall extinction-recolonisation reaction $W\rightarrow\emptyset\rightarrow M$ thus occurs at rate $1/(1/r_{{\rm ext},W}+1/r_{{\rm rec},M}^{\rm G})$. Here, the rate of $W$ deme extinction is $r_{{\rm ext},W}=(\Omega_{\text{occ}}-i)/\tau_{E}$ and $\tau_{E}$ is given by Eq. (7). We proceed similarly for the extinction of an $M$ deme and its recolonisation into a $W$ site according to $M\rightarrow\emptyset\rightarrow W$. Taking also into account the rate of invasion (see Eq. (11)) the number of $M$ demes $i$ on a regular graph $\rm G$} varies according to the transition rates
@@
It is important to note that the $\widetilde{E}_{\rm G}(i)$ in these expressions explicitly represent the number of active edges between $M$ and $W$ on the metapopulation given $i$ mutant demes, taking into account the presence of extinct demes. Therefore, their expressions for a given graph structure $\rm G$ generally differ from those considered in the competition-dominated regime, denoted $E_{\rm G}(i)$. With these rates, we can solve the following first-step analysis equations for the probability $\phi_{{\rm int},i}^{\rm G}$ that the dynamical equilibrium comprising initially $i$ demes of type $M$ consists of occupied $M$ demes and extinct demes after a mean time $\theta_{{\rm int},i}^{\rm G}$